Status Report

NOAA GOES SATOPS Morning Report 26 Mar 2001

By SpaceRef Editor
March 26, 2001
Filed under ,


SOCC Relocation Update – The installation of modular operations
furniture for the launch control and engineering analysis rooms was completed by March 23.
Currently, 32 DEC 3100 workstations are being configured for installation into the
new launch control room by the first week in April.


No change in the status of the POES spacecraft.

NOAA-16 Recorder 5B testing continues.

NOAA-15 rev 14865 / F at 1945z on March 23: No HRPT in real-time,
appeared to be a Suitland Earth Station Problem. GE Americom replaced a bad modem
at 2200z; no further problems reported.

POES operations were nominal over the past 72 hours.


No change in the status of the DMSP spacecraft.
DMSP operations were nominal over the past 72 hours.


No change in the status of the GOES spacecraft.

All eclipse operations nominal.

At 0945z on March 24, a GOES-8 frame break was observed and
attributed to RFI (unknown).

Between 1248z and 1552z on March 25, numerous GOES-10 frame
breaks were observed due to RFI. Operations failed over from the 18M ‘B’
antenna to the HR1 antenna (which was clean).

To review the eclipse schedules for the GOES-8 and GOES-10
spacecraft, refer to: to identify images
canceled to satisfy eclipse or instrument keep-out-zone (KOZ) requirements.

GOES-M Simulations – The third series of GOES-M Simulations will
be held March 28-30, with the participation of SS/L engineers as well as local
personnel. The scripts to be simulated will be Launch, Apogee Motor Fire#1, Mag Boom/Solar
Array Deploy, and Dipole Estimation. There will be two sessions of health and
safety crew simulations for the Crew 1 personnel.

GOES operations were nominal over the past 72 hours.

Questions or comments to

SpaceRef staff editor.